OpenStudy (anonymous):

Find the magnitude and direction angle of: v=6i+3j round direction angle to nearest degree

OpenStudy (bahrom7893):

magnitude is just sqrt(6^2+3^2)

OpenStudy (anonymous):

for finding magnitude you must sequre root (6^2+3^2)

OpenStudy (anonymous):

\[ \left| V \right| = \sqrt{6^2+3^2}=\sqrt{36+9}\sqrt{45}\]

OpenStudy (bahrom7893):

The angle is: Tan(theta) = 3/6

OpenStudy (bahrom7893):

y component / x component

OpenStudy (anonymous):

\[\tan \theta=\frac{ 3 }{ 6 }\]

OpenStudy (bahrom7893):

tan(theta) = 1/2 theta = ArcTan(1/2)

OpenStudy (bahrom7893):

about 27 degrees

OpenStudy (anonymous):

26.56505118 lol ;)

OpenStudy (bahrom7893):

s/he asked us to round it.

OpenStudy (anonymous):

yeah was just playin ;)